### v1

@@ 6.1 $\mathbf{F}\lambda$ solver for equal and unknown distortion
Based on [10], the epipolar constraint with equal and unknown radial distortion can be written as
$\begin{matrix}&&[&x^{\prime}_{d}x_{d}&x^{\prime}_{d}y_{d}&x^{\prime}_{d}&y^{\prime}_{d}x_{d}&y^{\prime}_{d}y_{d}&y^{\prime}_{d}&x_{d}&y_{d}&1&]&\cdot&\mathbf{f}&+&\\ +&\lambda&[&0&0&x_{d}r^{2}&0&0&y^{\prime}_{d}r^{2}&x_{d}r^{\prime 2}&y_{d}r^{\prime}&r^{2}+r^{\prime 2}&]&\cdot&\mathbf{f}&+&\\ +&\lambda^{2}&[&0&0&0&0&0&0&0&0&r^{2}r^{\prime 2}&]&\cdot&\mathbf{f}&=&0\enspace,\end{matrix}$ (3)
@@
$(\mathbf{A}_{0}+\lambda\mathbf{A}_{1}+\lambda^{2}\mathbf{A}_{2})\mathbf{f}=\mathbf{0}\enspace.$ (4)
For 9 point correspondences in the 9pt $\mathbf{F}\lambda$ solver, equation (4) defines a polynomial eigenvalue problem that can be solved by computing the eigenvalues of a $18\times 18$ matrix. In [10], it was shown how the number of solutions of (4) can be reduced from 18 to 10 by transforming the problem to an eigenvalue problem of size $10\times 10$. However, in [10] it was also noted that 4 of these 10 solutions have been found to be imaginary. In our case, we show that the 4 imaginary solutions can be directly removed and we only need to find the eigenvalues of a $6\times 6$ matrix. Since matrix $\mathbf{A}_{2}$ is singular while $\mathbf{A}_{0}$ is full-rank, we first let $\sigma=1/\lambda$. Then (4) can be written as
$(\mathbf{A}_{2}+\sigma\mathbf{A}_{1}+\sigma^{2}\mathbf{A}_{0})\mathbf{f}=\mathbf{0}\enspace.$ (5)
@@
$\mathbf{B}=\begin{bmatrix}\mathbf{0}&{\mathbf{I}}\\ -{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{2}&-{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{1}\end{bmatrix}\enspace.$ (6)
There are 8 zero columns in $\mathbf{A}_{2}$, and 4 zero columns in $\mathbf{A}_{1}$. To solve this problem efficiently, we use the technique from [18]: the zero columns in $-{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{2}$ and $-{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{1}$ can be removed together with the corresponding rows. In this case, the size of the matrix $\mathbf{B}$ is reduced to $6\times 6$, and we only need to find the eigenvalues of a $6\times 6$ matrix. Note that in the solver, we directly construct the reduced $6\times 6$ matrix and avoid computations on the matrix (6)
For the non-minimal case, i.e., the case where the number of point correspondences is larger than 9, $-{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{2}$ and $-{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{1}$ are solved using linear least squares (which can be efficiently solved using the ColPivHouseholderQR function in the Eigen library [11]).
@@ 6.2 $\mathbf{F}\lambda_{1}\lambda_{2}$ solver for different unknown distortions
Similarly to the equal distortion case, here we propose minor modifications to the 12pt solver for two cameras with different distortions. The original 12pt solver was proposed in [19]. Note that in the following equations we use $\lambda$ and $\lambda^{\prime}$, similarly to Eq. (1) in the main paper, and not $\lambda_{1}$ and $\lambda_{2}$ as used in the names of the solvers.
As suggested in[19], we consider $\lambda^{\prime}$ as a hidden variable, i.e., we ”hide” it into coefficient matrices. Then the epipolar constraint can be written as
$\begin{matrix}&&[&x^{\prime}_{d}x_{d}&x^{\prime}_{d}y_{d}&x^{\prime}_{d}&y^{\prime}_{d}x_{d}&y^{\prime}_{d}y_{d}&y^{\prime}_{d}&x_{d}&y_{d}&1&x^{\prime}_{d}r^{2}&y^{\prime}_{d}r^{2}&r^{2}&]&\cdot&\tilde{\mathbf{f}}&+&\\ +&\lambda^{\prime}&[&0&0&0&0&0&0&x_{d}r^{\prime 2}&y_{d}r^{\prime 2}&r^{\prime 2}&0&0&r^{2}r^{\prime 2}&]&\cdot&\tilde{\mathbf{f}}&=&0\enspace,\end{matrix}$ (7)
where $\tilde{\mathbf{f}}$ is a $12\times 1$ vector of the form $\tilde{\mathbf{f}}=[\mathbf{f};\lambda f_{3};\lambda f_{6};\lambda f_{9}]$, $\mathbf{f}$ is a vector containing the entries of the fundamental matrix, and $f_{i}$ is the $i$th entry of the vector $\mathbf{f}$ By stacking the rows of $n$ correspondences, (7) can be written as
$(\mathbf{C}_{0}+\lambda^{\prime}\mathbf{C}_{1})\tilde{\mathbf{f}}=\mathbf{0}\enspace.$ (8)
@@
$(\mathbf{C}_{1}+\sigma^{\prime}\mathbf{C}_{0})\tilde{\mathbf{f}}=\mathbf{0}\enspace.$ (9)
@@
$\mathbf{D}=-{\mathbf{C}}_{0}^{-1}{\mathbf{C}}_{1}\enspace.$ (10)
As noted in [19], there are eight infinite eigenvalues, and at most four finite real solutions to this problem. Here we show that these infinite eigenvalues can be removed in advance.
Since there are 8 zero columns in $\mathbf{C}_{1}$, based on the method proposed in [18], these zero columns can be removed together with the corresponding rows. In this case, we only need to find the eigenvalues of a $4\times 4$ matrix.
Similarly to the equal distortion case, for more than 12 point correspondences, $-{\mathbf{C}}_{0}^{-1}{\mathbf{C}}_{1}$ is solved using linear least squares.
@@ 7 Inverse of the One-parameter Division Model
$u(\mathbf{x},\lambda)=[x_{d},y_{d},1+\lambda(x_{d}^{2}+y_{d}^{2})]^{\top}\enspace.$ (11)
@@
$x_{u}=\frac{x_{d}}{1+\lambda(x_{d}^{2}+y_{d}^{2})}\enspace,\ y_{u}=\frac{y_{d}}{1+\lambda(x_{d}^{2}+y_{d}^{2})}\enspace.$ (12)
@@
$x_{u}^{2}+y_{u}^{2}=\frac{x_{d}^{2}+y_{d}^{2}}{(1+\lambda(x_{d}^{2}+y_{d}^{2}))^{2}}\enspace.$ (13)
@@
$r_{u}=\frac{r_{d}}{1+\lambda r_{d}^{2}}\enspace,$ (14)
@@
$r_{u}=\sqrt{x_{u}^{2}+y_{u}^{2}}\text{ and }\ r_{d}=\sqrt{x_{d}^{2}+y_{d}^{2}}\enspace.$ (15)
@@
$r_{d}=\frac{1-\sqrt{1-4\lambda r_{u}^{2}}}{2\lambda r_{u}}\enspace.$ (17)
@@
$\begin{split}x_{d}&=x_{u}(1+\lambda r_{d}^{2})\enspace,\\ y_{d}&=y_{u}(1+\lambda r_{d}^{2})\enspace,\end{split}$ (18)
@@ 8 Benchmark Details
We created a new benchmark consisting two scenes, ROTUNDA and CATHEDRAL, for measuring relative pose accuracy under radial distortion. For both scenes, we build upon existing images: For the ROTUNDA scene, we use 62 GoPro images, kindly provided by the authors of [20]. For the CATHEDRAL scene, we use 655 GoPro images, kindly provided by the authors of [37]. For both scenes, we recorded additional images with different cameras. To obtain ground truth poses and camera intrinsics including radial distortion for both the original and the newly recorded images, we used the RealityCapture software [1]. We configured RealityCapture to estimate the undistortion parameters for the images using the one-parameter division model that we use in all of our experiments, i.e., RealityCapture directly provides ground truth estimates for the undistortion parameters. We enforced that images taken by the same camera (with the same field of view) share the same intrinsic camera parameters. In the following, we briefly describe both scenes.
The ROTUNDA scene contains 157 outdoor images of a historical building captured by two mobile phone cameras (95 images in total) and one GoPro camera (62 images, provided by the authors of [20]). The GoPro images were captured using two different settings that affect the field-of-view and image distortion. Overall, images were taken with four different $\lambda$ values: $\left\{-1.50,-0.81,0.02,0.09\right\}$ (ground truth values provided by RealityCapture after normalization). Fig. 3 visualizes the ROTUNDA scene by showing a textured mesh of the scene together with the camera poses of the recorded images. Figures 4 and 5 show example images from the ROTUNDA scene.
The CATHEDRAL scene contains 2,734 outdoor images of a historical cathedral, captured by two mobile phone cameras (708 images in total), one GoPro camera (655 images, provided by the authors of [37]), and an Insta360 Ace Pro camera (1,358). Most of the images were extracted from videos captured while walking around the building. In addition, we are using 13 images from Flickr. The dataset contains images from cameras with 42 different $\lambda$ parameters. Their distribution is shown in Fig. 4 along with example images for both scenes. Fig. 6 shows more example images from the CATHEDRAL scene. Fig. 7 visualizes the CATHEDRAL scene by showing a colored mesh of the scene together with the camera poses of the recorded images.
Figure 8: Example image of the Courtyard scene from the ETH3D dataset, w.r.t. synthetically added distortions. We use undistortion parameters $\{0,-0.3,-0.6,-0.9,-1.2,-1.8\}$ and compute the distorted images using the inverse of the one-parameter division undistortion model. The corresponding images are shown in left-to-right and top-to-bottom order. For the parameters, we consider the one-parameter division model and $[-0.5,0.5]$ coordinate normalization.
@@ 9 Experiments
In order to provide a more intuitive understanding of the range of distortion used in this work, Fig. 8 visualizes the impact of various undistortion parameters on an image. As in the main paper, the parameters are defined for normalized coordinates in the range $[-0.5,0.5]^{2}$.
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 15.77 7.67 0.19 0.33 0.50 0.69 0.72 1160.02
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 27.79 18.45 0.08 0.16 0.29 0.65 0.62 1173.91
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 20.30 6.83 0.30 0.42 0.57 0.69 0.72 1112.99
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 00 16.16 3.67 0.42 0.54 0.66 0.34 0.15 1164.82
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 16.27 3.74 0.42 0.54 0.67 0.35 0.17 1164.85
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 35.34 17.58 0.13 0.21 0.34 0.65 0.68 1112.10
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 30.33 13.30 0.19 0.28 0.41 0.52 0.42 1166.39
@@
For experiments in the main paper, we use SuperPoint [8] features extracted on images without resizing. We only kept at most 2048 features and matched them with LightGlue (LG) [26]. We only considered images with sufficient overlap quantified by the co-visibility constraint proposed in [15]. We also only kept image pairs, which resulted in at least 20 matches.
We also used the same strategy with SIFT [27] local features. Note that due to the criterion of at least 20 matches for each image pair, the sets of image pairs for all datasets differ slightly. We report results for SIFT+LG for both GC-RANSAC and Poselib frameworks on the CATHEDRAL scene (see Tab. 3). We used 10,000 randomly sampled image pairs for both $\lambda_{1}=\lambda_{2}$ and $\lambda_{1}\neq\lambda_{2}$.
@@
[…]
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 28.38 16.58 0.09 0.18 0.31 0.87 0.85 1520.07
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 34.11 23.24 0.05 0.10 0.22 0.86 0.86 1870.98
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 16.89 3.66 0.30 0.47 0.61 0.15 0.15 1280.81
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 20.49 5.68 0.21 0.37 0.52 0.15 0.15 1343.89
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 9pt $\mathbf{F}\lambda$ $-0.9$ 17.67 13.96 0.31 0.46 0.58 0.25 0.14 1230.37
7pt $\mathbf{F}$ 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ $-0.9$ 19.48 16.30 0.20 0.36 0.52 0.30 0.27 1291.12
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ $-0.9$ 26.80 12.35 0.11 0.23 0.38 0.33 0.31 1560.78
7pt $\mathbf{F}$ 8pt $\mathbf{F}$ $-0.9$ 26.96 12.10 0.11 0.22 0.38 0.33 0.31 1470.90

### v2

@@ 6.1 $\mathbf{F}\lambda$ solver for equal and unknown distortion
Based on [fitzgibbon2001simultaneous], the epipolar constraint with equal and unknown radial distortion can be written as
[…]
@@
$(\mathbf{A}_{0}+\lambda\mathbf{A}_{1}+\lambda^{2}\mathbf{A}_{2})\mathbf{f}=\mathbf{0}\kern 5.0pt.$ (4)
For 9 point correspondences in the 9pt $\mathbf{F}\lambda$ solver, equation (4) defines a polynomial eigenvalue problem that can be solved by computing the eigenvalues of a $18\times 18$ matrix. In [fitzgibbon2001simultaneous], it was shown how the number of solutions of (4) can be reduced from 18 to 10 by transforming the problem to an eigenvalue problem of size $10\times 10$. However, in [fitzgibbon2001simultaneous] it was also noted that 4 of these 10 solutions have been found to be imaginary. In our case, we show that the 4 imaginary solutions can be directly removed and we only need to find the eigenvalues of a $6\times 6$ matrix. Since matrix $\mathbf{A}_{2}$ is singular while $\mathbf{A}_{0}$ is full-rank, we first let $\sigma=1/\lambda$. Then (4) can be written as
$(\mathbf{A}_{2}+\sigma\mathbf{A}_{1}+\sigma^{2}\mathbf{A}_{0})\mathbf{f}=\mathbf{0}\kern 5.0pt.$ (5)
@@
$\mathbf{B}=\begin{bmatrix}\mathbf{0}&{\mathbf{I}}\\ -{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{2}&-{\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{1}\end{bmatrix}\kern 5.0pt$. (6)
There are 8 zero columns in $\mathbf{A}_{2}$, and 4 zero columns in $\mathbf{A}_{1}$. To solve this problem efficiently, we use the technique from [kukelova2012polynomial]: the zero columns in ${\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{2}$ and ${\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{1}$ can be removed together with the corresponding rows. In this case, the size of the matrix $\mathbf{B}$ is reduced to $6\times 6$, and we only need to find the eigenvalues of a $6\times 6$ matrix. Note that in the solver, we directly construct the reduced $6\times 6$ matrix and avoid computations on the matrix (6)
For the non-minimal case, i.e., the case where the number of point correspondences is larger than 9, ${\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{2}$ and ${\mathbf{A}}_{0}^{-1}{\mathbf{A}}_{1}$ are solved using linear least squares (which can be efficiently solved using the ColPivHouseholderQR function in the Eigen library [eigenweb]).
@@ 6.2 $\mathbf{F}\lambda_{1}\lambda_{2}$ solver for different unknown distortions
Similarly to the equal distortion case, here we propose minor modifications to the 12pt solver for two cameras with different distortions. The original 12pt solver was proposed in [kukelova2010fast]. Note that in the following equations we use $\lambda$ and $\lambda^{\prime}$, similarly to Eq. (1) in the main paper, and not $\lambda_{1}$ and $\lambda_{2}$ as used in the names of the solvers.
As suggested in[kukelova2010fast], we consider $\lambda^{\prime}$ as a hidden variable, i.e., we ”hide” it into coefficient matrices. Then the epipolar constraint can be written as
[…]
where $\tilde{\mathbf{f}}$ is a $12\times 1$ vector of the form $\tilde{\mathbf{f}}=[\mathbf{f};\lambda f_{3};\lambda f_{6};\lambda f_{9}]$, $\mathbf{f}$ is a vector containing the entries of the fundamental matrix, and $f_{i}$ is the $i$th entry of the vector $\mathbf{f}$ . By stacking the rows of $n$ correspondences, (7) can be written as
$(\mathbf{C}_{0}+\lambda^{\prime}\mathbf{C}_{1})\tilde{\mathbf{f}}=\mathbf{0}\kern 5.0pt.$ (8)
@@
$(\mathbf{C}_{1}+\sigma^{\prime}\mathbf{C}_{0})\tilde{\mathbf{f}}=\mathbf{0}\kern 5.0pt.$ (9)
@@
$\mathbf{D}=-{\mathbf{C}}_{0}^{-1}{\mathbf{C}}_{1}\kern 5.0pt.$ (10)
As noted in [kukelova2010fast], there are eight infinite eigenvalues, and at most four finite real solutions to this problem. Here we show that these infinite eigenvalues can be removed in advance.
Since there are 8 zero columns in $\mathbf{C}_{1}$, based on the method proposed in [kukelova2012polynomial], these zero columns can be removed together with the corresponding rows. In this case, we only need to find the eigenvalues of a $4\times 4$ matrix.
Similarly to the equal distortion case, for more than 12 point correspondences, ${\mathbf{C}}_{0}^{-1}{\mathbf{C}}_{1}$ is solved using linear least squares.
@@ 7 Inverse of the One-parameter Division Model
$u(\mathbf{x},\lambda)=[x_{d},y_{d},1+\lambda(x_{d}^{2}+y_{d}^{2})]^{\top}\kern 5.0pt.$ (11)
@@
$x_{u}=\frac{x_{d}}{1+\lambda(x_{d}^{2}+y_{d}^{2})}\kern 5.0pt,\ y_{u}=\frac{y_{d}}{1+\lambda(x_{d}^{2}+y_{d}^{2})}\kern 5.0pt.$ (12)
@@
$x_{u}^{2}+y_{u}^{2}=\frac{x_{d}^{2}+y_{d}^{2}}{(1+\lambda(x_{d}^{2}+y_{d}^{2}))^{2}}\kern 5.0pt.$ (13)
@@
$r_{u}=\frac{r_{d}}{1+\lambda r_{d}^{2}}\kern 5.0pt,$ (14)
@@
$r_{u}=\sqrt{x_{u}^{2}+y_{u}^{2}}\text{ and }\ r_{d}=\sqrt{x_{d}^{2}+y_{d}^{2}}\kern 5.0pt.$ (15)
@@
$r_{d}=\frac{1-\sqrt{1-4\lambda r_{u}^{2}}}{2\lambda r_{u}}\kern 5.0pt.$ (17)
@@
$\begin{split}x_{d}&=x_{u}(1+\lambda r_{d}^{2})\kern 5.0pt,\\ y_{d}&=y_{u}(1+\lambda r_{d}^{2})\kern 5.0pt,\end{split}$ (18)
@@ 8 Benchmark Details
We created a new benchmark consisting two scenes, ROTUNDA and CATHEDRAL, for measuring relative pose accuracy under radial distortion. For both scenes, we build upon existing images: For the ROTUNDA scene, we use 62 GoPro images, kindly provided by the authors of [kukelova2015efficient]. For the CATHEDRAL scene, we use 655 GoPro images, kindly provided by the authors of [Sattler_2019_CVPR]. For both scenes, we recorded additional images with different cameras. To obtain ground truth poses and camera intrinsics including radial distortion for both the original and the newly recorded images, we used the RealityCapture software [RealityCapture]. We configured RealityCapture to estimate the undistortion parameters for the images using the one-parameter division model that we use in all of our experiments, i.e., RealityCapture directly provides ground truth estimates for the undistortion parameters. We enforced that images taken by the same camera (with the same field of view) share the same intrinsic camera parameters. In the following, we briefly describe both scenes.
The ROTUNDA scene contains 157 outdoor images of a historical building captured by two mobile phone cameras (95 images in total) and one GoPro camera (62 images, provided by the authors of [kukelova2015efficient]). The GoPro images were captured using two different settings that affect the field-of-view and image distortion. Overall, images were taken with four different $\lambda$ values: $\left\{-1.50,-0.81,0.02,0.09\right\}$ (ground truth values provided by RealityCapture after normalization). Fig. 3 visualizes the ROTUNDA scene by showing a textured mesh of the scene together with the camera poses of the recorded images. Figures 4 and 5 show example images from the ROTUNDA scene.
The CATHEDRAL scene contains 2,734 outdoor images of a historical cathedral, captured by two mobile phone cameras (708 images in total), one GoPro camera (655 images, provided by the authors of [Sattler_2019_CVPR]), and an Insta360 Ace Pro camera (1,358). Most of the images were extracted from videos captured while walking around the building. In addition, we are using 13 images from Flickr. The dataset contains images from cameras with 42 different $\lambda$ parameters. Their distribution is shown in Fig. 4 along with example images for both scenes. Fig. 6 shows more example images from the CATHEDRAL scene. Fig. 7 visualizes the CATHEDRAL scene by showing a colored mesh of the scene together with the camera poses of the recorded images.
Figure 8: Example image of the Courtyard scene from the ETH3D dataset, w.r.t. synthetically added distortions. We use undistortion parameters $\{0,-0.3,-0.6,-0.9,-1.2,-1.8\}$ and compute the distorted images using the inverse of the one-parameter division undistortion model. The corresponding images are shown in left-to-right and top-to-bottom order. For the parameters, we consider the one-parameter division model and [0.5,0.5][-0.5,0.5] coordinate normalization.
@@ 9 Experiments
In order to provide a more intuitive understanding of the range of distortion used in this work, Fig. 8 visualizes the impact of various undistortion parameters on an image. As in the main paper, the parameters are defined for normalized coordinates in the range $0.5,0.5]^{2}$.
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 15.77 7.67 0.19 0.33 0.50 0.69 0.72 1160.02
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 27.79 18.45 0.08 0.16 0.29 0.65 0.62 1173.91
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 20.30 6.83 0.30 0.42 0.57 0.69 0.72 1112.99
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0 16.16 3.67 0.42 0.54 0.66 0.34 0.15 1164.82
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 16.27 3.74 0.42 0.54 0.67 0.35 0.17 1164.85
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 35.34 17.58 0.13 0.21 0.34 0.65 0.68 1112.10
@@
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 30.33 13.30 0.19 0.28 0.41 0.52 0.42 1166.39
@@
For experiments in the main paper, we use SuperPoint [detone2018superpoint] features extracted on images without resizing. We only kept at most 2048 features and matched them with LightGlue (LG) [lindenberger2023lightglue]. We only considered images with sufficient overlap quantified by the co-visibility constraint proposed in [IMC2020]. We also only kept image pairs, which resulted in at least 20 matches.
We also used the same strategy with SIFT [Lowe_sift] local features. Note that due to the criterion of at least 20 matches for each image pair, the sets of image pairs for all datasets differ slightly. We report results for SIFT+LG for both GC-RANSAC and Poselib frameworks on the CATHEDRAL scene (see Tab. 3). We used 10,000 randomly sampled image pairs for both $\lambda_{1}=\lambda_{2}$ and $\lambda_{1}\neq\lambda_{2}$.
@@
Table 4: Results for {($\{$(Scenario A) Wild, (Scenario B) Small distortion, (Scenario C) Visible distortion}\} on 12 scenes from the ETH3D dataset, using the GC-RANSAC framework and SP+LG matches.
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 28.38 16.58 0.09 0.18 0.31 0.87 0.85 1520.07
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 34.11 23.24 0.05 0.10 0.22 0.86 0.86 1870.98
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 16.89 3.66 0.30 0.47 0.61 0.15 0.15 1280.81
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 20.49 5.68 0.21 0.37 0.52 0.15 0.15 1343.89
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 9pt $\mathbf{F}\lambda$ 0.9-0.9 17.67 13.96 0.31 0.46 0.58 0.25 0.14 1230.37
7pt $\mathbf{F}$ 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ 0.9-0.9 19.48 16.30 0.20 0.36 0.52 0.30 0.27 1291.12
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0.9-0.9 26.80 12.35 0.11 0.23 0.38 0.33 0.31 1560.78
7pt $\mathbf{F}$ 8pt $\mathbf{F}$ 0.9-0.9 26.96 12.10 0.11 0.22 0.38 0.33 0.31 1470.90